# EgoAgent: A Joint Predictive Agent Model in Egocentric Worlds

### v2

@@ Abstract
This paper addresses the task of learning an agent model behaving like humans, which can jointly perceive, predict, and act in egocentric worlds. Previous methods usually train separate models for these three abilities, which prevents them from learning from each other. In this paper, we propose a joint predictive agent model, named EgoAgent, that simultaneously learns to represent the world, predict future states, and take reasonable actions within a single transformer. EgoAgent introduces two innovations to learn from the causal and temporally intertwined nature of these abilities: (1) Interleaved sequential modeling of states and actions with the causal attention mechanism, and (2) A joint embedding-action-prediction architecture featuring temporal asymmetric predictor-observer branches. Integrating these designs based on JEPA [29], EgoAgent unifies these capabilities in a cohesive learning framework. Comprehensive evaluations of EgoAgent on representative tasks such as image classification, egocentric future state prediction, and 3D human motion prediction tasks demonstrate the superiority of our method. The code and trained model will be released for reproducibility.
Figure 1: (a) Humans learn from the world by continuously perceiving the egocentric world, predicting future states, and taking actions to achieve their goals. (b) Inspired by the human learning process, we proposed EgoAgent, a joint predictive agent model in egocentric worlds that can learn to represent, predict, and act from egocentric video and 3D human skeleton sequences.
@@ 1 Introduction
Cognition is embodied; it arises from bodily interactions with the world,” as stated by Linda B. Smith [^1] Linda Brawn Smith is a cognitive scientist recognized for her work in developmental psychology and cognitive science, with over 52,000 citations. She won the David E. Rumelhart Prize for theoretical contributions to cognitive science and is a member of both the National Academy of Sciences and the American Academy of Arts and Science. in [43]. This perspective emphasizes that cognition emerges from continuous, dynamic interactions between the brain, body, and environment, rather than relying on symbolic systems detached from real-world experiences [42, 8, 39]. Through these ongoing interactions, humans naturally acquire foundational abilities such as visual perception, predictive understanding of world dynamics, and action anticipation. As highlighted by Yann LeCun [29], replicating these abilities in artificial systems remains one major challenge in AI research. Motivated by this human learning process, we propose an agent model that jointly learns these abilities from egocentric RGB observations and their responsive actions. Such abilities are essential for applications in robotics, augmented reality, and gaming where agents must understand what is happening in the world, anticipate what will happen next, and decide what actions to take.
@@
Existing approaches typically address these capabilities in isolated tasks: (1) Visual representation learning [10, 26, 46], which encodes high-level representations of human observations of the world; (2) World models [36, 52], which learn predictive representations of world state transitions conditioned on actions. (3) Action prediction [35, 9], which forecasts future human actions based on prior motion. However, research in cognitive science, particularly the famous Common Coding Theory [38, 8], claims that perception and action are not separate processes but deeply intertwined and mutually reinforcing, operating within a shared representational space. Therefore, modeling these abilities into a unified framework can be not only mutually beneficial but also fundamentally feasible for agent models to achieve human-like understanding in egocentric worlds.
@@
While combining these abilities offers clear benefits, integrating them into one model remains challenging. Humans interact with the world through an interleaved cycle of perceiving egocentric observations and taking actions, creating close interdependencies between observations and actions in time and causality. To mimic and learn from this behavior, researchers must model the causal and temporally interlaced nature of observations and actions within the network, and design effective supervision signals to capture their dependencies, which is an unsolved problem.
@@
In response, we propose EgoAgent, a joint predictive agent model that simultaneously learns to represent world observations, predict future states, and act based on learned representations (Fig. 1). To model and learn from the intertwined nature of these abilities, EgoAgent introduces two key innovations: (1) Formulating human-world interactions as interleaved sequences of “state-action-state-action” tokens processed by the causal attention mechanism. This design explicitly captures the interlaced structure of egocentric observations and responsive actions across temporal and causal dimensions within a unified representational space. (2) A joint embedding-action-prediction (JEAP) architecture with temporal asymmetric predictor-observer branches, where the predictor branch leverages past information to predict future world states while the observer extracts the target world states from raw observations. By minimizing the prediction error in feature spaces, JEAP is tailored to learn the temporal transitions of the world, thereby integrating perception, action, and prediction into a cohesive learning framework.
@@
Our contributions are as follows: (1) We develop the first agent model that can simultaneously learn to represent egocentric observations, predict future states, and generate actions in a unified representational space. (2) We propose a JEAP architecture with temporal asymmetric branches to learn both semantic understanding and temporal transitions of the dynamic multi-modal egocentric world. (3) We evaluate our EgoAgent on egocentric video-action datasets and show its superior abilities in visual representation, world states prediction, and 3D human motion prediction tasks.
@@ 2 Related Work
Figure 2: The overall framework of EgoAgent. EgoAgent adopts the Joint Embedding-Action-Prediction Architecture to project input egocentric video frames $I$ and 3D human poses $A$ into image tokens $i$ and action tokens $a$. Learnable action query token $q_{a}$ and future state query token $q_{s}$ are appended after the image and action tokens to stimulate EgoAgent’s predictor $\Theta_{pred}$ to predict the next action $A^{\prime}$ and world state $S^{\prime}$. The action prediction loss $\mathcal{L}_{act}$ is used to supervise the predicted actions. Since EgoAgent predicts world states $S^{\prime}$ in a semantic feature space, the target world state $S$ is obtained from a momentum observer $\Theta_{obs}$. Features from the predictor and observer branches are aligned through a state prediction loss $\mathcal{L}_{pred}$. When 3D poses are not paired with video frames, EgoAgent randomly crops different views from the video and learns representative features only with image tokens $i$ and state query tokens $q_{s}$ by a representation loss $\mathcal{L}_{rep}$. For easier understanding, we only illustrate the supervisions on predicting the scene state and action at time step $t$.
@@
World Models and Agent Models. Ha and Schmidhuber [24] defined the world model as a predictive representation of the environment, while the agent model utilizes this learned representation to make decisions and take actions. Recently, generative world models [52, 2, 49] have gained significant attention, often employing an autoencoder framework to simulate transitions between video frames. These models have demonstrated success in applications such as autonomous driving [47], robotics [23], and game control [5]. For instance, GAIA-1 [27] maps multi-modal driving signals into discrete tokens, and processes them using an autoregressive transformer for unified representation learning. In contrast, predictive world models [3, 7, 18] focus on learning state transitions through the joint embedding predictive architecture (JEPA) [29] without generating videos explicitly. For example, MC-JEPA [6] incorporates self-supervised learning with optical flow estimation to learn content features and motion dynamics. In this work, we extend the JEPA world model into an agent model in egocentric worlds, which not only predicts world states based on human actions but also predicts actions based on learned world state transitions.
@@
Egocentric Visual Representation Learning. Traditional visual representation methods learn the features on large-scale curated image datasets, such as ImageNet [15], using either supervised [40, 28, 46] or self-supervised methods [25, 10, 12]. However, these datasets primarily consist of third-person perspectives, which do not fully align with how humans perceive the world. To bridge this gap, researchers have turned to egocentric videos, which offer a more natural and immersive viewpoint for learning visual representations. Models like R3M [37] and VIP [31] leverage the temporal relationships in egocentric videos to learn generalizable representations from Ego4D [19] for robotics applications. Recently, DoRA [44] learns object-level representations by tracking movements on an egocentric video dataset, WalkingTours. In this paper, we propose an agent model that not only learns robust representations from egocentric videos but also predicts future states and actions.
@@ 3 Method
Our goal is to develop an agent model that can represent world states, predict future states, and generate the following actions by simulating how humans interact with the world through egocentric vision and 3D skeletal motion. The overview of our framework is presented in Fig. 2. We first introduce our joint embedding-action-prediction architecture in Sec. 3.1. We then introduce how EgoAgent can be trained in Sec. 3.2. Finally, we demonstrate the application of EgoAgent in the three tasks in Sec. 3.3.
@@ 3.1 Joint Embedding-Action-Prediction Architecture for EgoAgent
Given egocentric observations $V=\{I_{1},I_{2},...I_{t}\}$ and history human actions $A=\{A_{1},A_{2},...A_{t}\}$, our agent model aims to predict the future world state $S_{t+1}$ associated with the next observation $I_{t+1}$ while simultaneously generate plausible human action $A_{t+1}$ by understanding world states $S_{[0:t]}$. The world states are encoded as semantic embeddings, whereas human actions are represented as 3D poses.
@@
Joint prediction pipeline. Our pipeline begins by encoding the egocentric video $V$ and human action sequence $A$ into semantic features. For egocentric video, we follow the approach of ViT [16], applying a convolutional layer to each video frame $I_{t}$ to produce a feature map. This map is then subdivided and flattened to obtain the final image feature $i_{t}$. For human actions, we utilize a convolutional layer followed by layer normalization (LN) and a Gaussian error linear unit (GeLU) activation to map the 3D skeletons $A_{t}$ into action features $a_{t}$. These feature vectors are then processed by a causal transformer network to generate the target outputs. As illustrated in Fig. 2, the input sequence is structured as a sequence of image, action, and query tokens.
@@
At each time step $t$, we assemble a structured sequence of tokens, including the image token $i_{t}$, an action query token $q_{a}$, the action token $a_{t}$, and a future state query token $q_{s}$, where the query tokens are learnable embeddings. Enabled by the causal attention mechanism of the transformer, $q_{s}$ prompts the model to incorporate both previous and current image tokens $i_{[0:t]}$ and action tokens $a_{[0:t]}$ for predicting $s_{t+1}$. Notably, we insert $q_{a}$ right after the image tokens and before the action tokens to prevent the model from accessing the future action $a_{t}$ during prediction. This ensures that the model relies solely on past observations $i_{[0:t]}$ and past actions $a_{[0:t-1]}$ when generating future actions. Subsequently, the transformer network processes these tokens and outputs the action embedding $a_{t}^{{}^{\prime}}$ and future state embedding $s_{t+1}^{{}^{\prime}}$ in response to the query tokens. Finally, we utilize separate MLP networks to map these embeddings to the predicted actions $A_{t}^{{}^{\prime}}$ and future world states $S_{t+1}^{{}^{\prime}}$.
@@
JEPA with temporal asymmetric branches. Unlike actions, which can be directly obtained and supervised, world states must be learned from observations while also be used as supervision signals for state prediction. To achieve this, we propose a temporal asymmetric predictor-observer architecture based on JEPA [29, 18], a framework that combines representation learning with predictive learning. Despite of the predictor branch executing the joint prediction pipeline, we introduce an observer branch that only processes image inputs by extracting current-frame features for representation learning supervision and providing next-frame features for future state supervision alongside the aforementioned predictor branch. Specifically, the observer branch takes image tokens $i_{t}$ and state query tokens $q_{s}$ as input, and outputs current state embeddings $s_{t}$, which is fed into the same MLP network used for image embeddings in the observer branch to obtain world states $S_{t}$.
@@
Learning in semantic feature space. Notably, unlike previous methods [4, 52] that rely on pretrained reconstruction-based tokenizers such as VQGAN [17] to convert images into discrete tokens, we employ learnable convolutional layers to project images into continuous semantic embeddings. Discrete tokens prioritize pixel-level reconstruction over high-level semantic information, whereas humans make predictions based on abstract concepts rather than by picturing at the pixel level. Therefore, our approach aligns more closely with human reasoning, enhancing the model’s performance in future state prediction.
@@ 3.2 Training EgoAgent
where $sg[\cdot]$ denotes the stop-gradient operation, $\mathcal{L}_{1}$ and $\mathcal{L}_{dino}$ represent the L1 loss and the DINO loss [10], respectively. Following common practices in self-supervised learning [10, 25], sg[$\cdot$] is applied to block the gradients from back-propagating to the teacher branch. The weights of the observer branch are updated in each iteration using an exponential moving average (EMA) of the predictor branch.
Self-supervision for learning powerful representations. When humans learn to interact with the environment, they first develop an understanding of the observed world and objects, which then aids them in predicting future states of the world and making appropriate responses. Inspired by this process, we introduce an additional self-supervised learning loss on EgoAgent to facilitate learning representative features from egocentric videos from scratch:
@@
Overall objective function. Given an input token sequence containing $t$ time steps, we have the overall objective function $\mathcal{L}$ defined as:
@@
where $\lambda_{rep}$, $\lambda_{pred}$, and $\lambda_{act}$ are the corresponding loss weights for representing the world, predicting future world states and actions, respectively.

### v3

@@ Abstract
Learning an agent model that behaves like humans—capable of jointly perceiving the environment, predicting the future, and taking actions from a first-person perspective—is a fundamental challenge in computer vision. Existing methods typically train separate models for these abilities, which fail to capture their intrinsic relationships and prevent them from learning from each other. Inspired by how humans learn through the perception-action loop, we propose EgoAgent, a unified agent model that simultaneously learns to represent, predict, and act within a single transformer. EgoAgent explicitly models the causal and temporal dependencies among these abilities by formulating the task as an interleaved sequence of states and actions. It further introduces a joint embedding–action–prediction architecture with temporally asymmetric predictor and observer branches, enabling synergistic optimization across all three capabilities. Comprehensive evaluations of EgoAgent on representative tasks such as image classification, egocentric future state prediction, and 3D human motion prediction demonstrate the superiority of our method. The code and trained models will be publicly available at https://github.com/zju3dv/EgoAgent.
Figure 1: (a) Humans learn from the world by continuously perceiving the egocentric world, predicting future states, and taking actions to achieve their goals. (b) Inspired by the human learning process, we propose EgoAgent, a joint predictive agent model in egocentric worlds that can learn to represent, predict, and act from egocentric videos and synchronized 3D human body motions.
@@ 1 Introduction
Cognition is embodied; it arises from bodily interactions with the world,” as stated by Linda B. Smith [^1] Linda Brawn Smith is a cognitive scientist recognized for her work in developmental psychology and cognitive science, with over 53,000 citations. She won the David E. Rumelhart Prize for theoretical contributions to cognitive science and is a member of both the National Academy of Sciences and the American Academy of Arts and Science. in [43]. This perspective emphasizes that cognition emerges from continuous, dynamic interactions between the brain, body, and environment, rather than relying on symbolic systems detached from real-world experiences [42, 8, 40]. Through these ongoing interactions, humans naturally acquire foundational abilities such as visual perception, predictive understanding of world dynamics, and action anticipation. Replicating these abilities in artificial systems remains one major challenge in AI research, as highlighted by Yann LeCun [30]. Motivated by this human learning process, we propose an agent model that jointly learns these abilities from egocentric RGB observations and their responsive actions. Such abilities are essential for applications in robotics, augmented reality, and gaming, where agents must understand what is happening in the world, anticipate what will happen next, and decide what actions to take.
@@
Existing approaches typically address these capabilities in isolated tasks: (1) Visual representation learning [10, 27, 46], which encodes high-level representations of human observations of the world; (2) World models [37, 51], which learn predictive representations of world state transitions conditioned on actions; (3) Action prediction [36, 9], which forecasts future human actions based on prior motion. However, research in cognitive science, particularly the famous Common Coding Theory [39, 8], claims that perception and action are not separate processes but deeply intertwined and mutually reinforcing, operating within a shared representational space. Therefore, modeling these abilities within a unified framework is not only mutually beneficial but also represents a more principled approach for agent models to achieve human-like understanding in egocentric worlds.
@@
While combining these abilities offers clear benefits, integrating them into one model remains challenging. Human interaction with the world involves a continuous loop of perceiving egocentric observations and taking 3D actions, creating complex interdependencies between observations and actions in time and causality. A successful model must therefore learn from this interleaved and tightly-coupled structure. The core difficulty, which remains an open question, lies in designing learning frameworks and supervision signals to capture these intricate connections.
@@
In response, we propose EgoAgent, a joint predictive agent model that simultaneously learns to represent world observations, predict future states, and act based on learned representations (Fig. 1). EgoAgent introduces two key innovations: (1) Formulating human-world interactions as interleaved sequences of “state-action-state-action” tokens processed by the causal attention mechanism. This design explicitly models both how observations trigger actions and how those actions, in turn, influence the subsequent states of the world. (2) A joint embedding-action-prediction (JEAP) architecture with temporally asymmetric predictor-observer branches. The predictor leverages past information to forecast future human actions and world states, while the observer extracts the target world states from raw observations. By minimizing the prediction error between these two branches in a shared semantic feature space, JEAP is tailored to learn both world dynamics and human policy based on high-level visual representations, thereby integrating the three tasks into a cohesive learning framework.
@@
Our contributions are as follows: (1) We develop the first agent model that can simultaneously learn to represent egocentric observations, predict future states, and generate 3D actions in a unified representational space. (2) We propose a JEAP architecture with temporally asymmetric branches to learn both semantic understanding and temporal transitions of the dynamic multi-modal egocentric world. (3) We evaluate our EgoAgent on egocentric video-action datasets and show its superior abilities in visual representation, world states prediction, and 3D human motion prediction tasks.
@@ 2 Related Work
Figure 2: The overall framework of EgoAgent. EgoAgent adopts the Joint Embedding-Action-Prediction Architecture to project input egocentric video frames $I_{t}$ and 3D human poses $A_{t}$ into image tokens $i_{t}$ and action tokens $a_{t}$. Learnable action query token $q_{a}$ and future state query token $q_{s}$ are appended after the image and action tokens to stimulate EgoAgent’s predictor $\Theta_{pred}$ to predict the next action $A^{\prime}_{t}$ and world state $S^{\prime}_{t+1}$. The action prediction loss $\mathcal{L}_{act}$ is used to supervise the predicted actions. Since EgoAgent predicts world states $S^{\prime}_{t+1}$ in a semantic feature space, the target world state $S_{t+1}$ is obtained from a momentum observer $\Theta_{obs}$. Features from the predictor and observer are aligned through a state prediction loss $\mathcal{L}_{pred}$. When 3D poses are not paired with video frames, EgoAgent randomly crops different views from the video and learns representative features only with image tokens $i_{t}$ and state query tokens $q_{s}$ by a self-supervised representation loss $\mathcal{L}_{rep}$. For easier understanding, we only illustrate the supervisions on predicting the state $S_{t}$ and action $A_{t}$.
@@
World Models and Agent Models. Ha and Schmidhuber [23] defined the world model as a predictive representation of the environment, while the agent model utilizes this learned representation to make decisions and take actions. Yann LeCun [30] further envisioned that an agent model can generate actions either reactively or through iterative optimization using the world model. Recently, generative world models [51, 2, 49] have demonstrated success in autonomous driving [47], robotics [22], and game control [5], mostly employing an encoder-decoder architecture to learn transitions between video frames. GAIA-1 [28] maps multi-modal driving signals into discrete tokens, and processes them using an autoregressive transformer. In contrast, JEPA world models [3, 7, 18] focus on learning state transitions through the joint embedding predictive architecture (JEPA) [30] without generating videos explicitly. MC-JEPA [6] incorporates self-supervised learning with optical flow estimation to learn content features and motion dynamics. In this work, we extend the JEPA world model into a reactive agent model, which not only predicts world states based on human actions but also predicts actions based on learned world state transitions.
@@
Egocentric Visual Representation Learning. Traditional visual representation learning methods are trained on large-scale image datasets, such as ImageNet [15], using either supervised [25, 29, 46] or self-supervised methods [26, 10, 12]. However, these datasets primarily consist of third-person perspectives, which do not fully align with how humans perceive the world. To bridge this gap, researchers have turned to egocentric videos, which offer a more natural and immersive viewpoint for learning visual representations. Models like R3M [38] and VIP [32] leverage the temporal relationships in egocentric videos to learn generalizable representations from Ego4D [19] for robotics applications. Recently, DoRA [44] learns object-level representations by tracking movements on an egocentric long video dataset called WalkingTours. In this paper, we propose an agent model that not only learns robust representations from egocentric videos but also predicts future states and actions.
@@ 3 Method
Our goal is to develop an agent model that can represent world states, predict future states, and generate future actions by simulating how humans interact with the world through egocentric vision and 3D skeletal motion. The overview of our framework is presented in Fig. 2. We first introduce our joint embedding-action-prediction architecture in Sec. 3.1. We then introduce how EgoAgent can be trained in Sec. 3.2. Finally, we demonstrate the application of EgoAgent in the three tasks in Sec. 3.3.
@@ 3.1 Joint Embedding-Action-Prediction Architecture for EgoAgent
Given egocentric observations $V=\{I_{1},I_{2},...I_{t}\}$ and human actions $A=\{A_{1},A_{2},...A_{t}\}$, our agent model aims to predict the future world state $S_{t+1}$ associated with the next observation $I_{t+1}$ while simultaneously generate plausible human action $A_{t+1}$. The world states are represented as semantic embeddings, whereas human actions are 3D poses.
@@
Interleaved joint prediction. EgoAgent begins by encoding the egocentric video $V$ and human action sequence $A$ into semantic embeddings. For video frames, we follow the approach of ViT [16], applying a convolutional layer to each video frame $I_{t}$ to produce a feature map. This map is then subdivided and flattened to obtain the image feature $i_{t}$. For human actions, we utilize a convolutional layer followed by layer normalization (LN) and a Gaussian error linear unit (GeLU) activation to map the 3D body pose $A_{t}$ into action feature $a_{t}$. These feature vectors are then processed by a causal transformer network.
@@
As illustrated in Fig. 2, at each time step $t$, we assemble a structured sequence of interleaved tokens, including the image token $i_{t}$, an action query token $q_{a}$, the action token $a_{t}$, and a future state query token $q_{s}$. The query tokens are learnable embeddings that serve as explicit prompts, directing the model to predict either the next action or a future world state. Notably, we insert $q_{a}$ right after the image tokens and $q_{s}$ after the action tokens to model the causal and temporally interlaced nature of perceiving observations and taking actions. Specifically, enabled by transformer’s causal attention mechanism, $q_{a}$ incorporates all preceding image tokens $i_{[0:t]}$ and action tokens $a_{[0:t-1]}$ to predict $a^{\prime}_{t}$. Similarly, $q_{s}$ integrates $i_{[0:t]}$ and the newly appended $a_{[0:t]}$ to predict $s^{\prime}_{t+1}$. We utilize separate MLPs to map these embeddings to 3D actions $A^{\prime}_{t}$ and world states $S^{\prime}_{t+1}$.
@@
JEPA with temporally asymmetric branches. Unlike actions, which can be directly obtained and supervised, world states must be learned from observations while also be used as supervision signals for state prediction. To achieve this, we propose a temporally asymmetric predictor-observer architecture based on JEPA [30, 18], a framework that combines representation learning with predictive learning. In addition to the predictor branch executing the interleaved joint prediction, we introduce an observer branch that only processes image inputs by (1) extracting current-frame features for self-supervised representation learning and (2) providing next-frame features for future state supervision alongside the predictor. Specifically, the observer branch takes image tokens $i_{t}$ and state query tokens $q_{s}$ as input, and outputs current state embeddings $s_{t}$, which are then fed into the same MLP for obtaining world states $S_{t}$ from image embeddings in the predictor. A key advantage of our query-based design is that it effectively decouples the shared state/representation components from the predictor’s action-specific ones. This prevents conflicting gradient updates, allowing the observer’s parameters to be updated as an exponential moving average (EMA) of the predictor’s, despite the observer not processing the action modality.
@@
Learning in semantic feature space. Notably, unlike previous methods [4, 51] that rely on pretrained reconstruction-based tokenizers, such as VQGAN [17], to convert images into discrete tokens, we employ learnable convolutional layers to project images into continuous semantic embeddings. Discrete tokens prioritize pixel-level reconstruction over high-level semantic information, whereas humans make predictions based on abstract concepts rather than pixels. Therefore, our approach aligns more closely with human reasoning and significantly enhances the model’s performance in future state prediction (see Sec. 4.2).
@@ 3.2 Training EgoAgent
where $sg[\cdot]$ denotes the stop-gradient operation, $\mathcal{L}_{1}$ and $\mathcal{L}_{dino}$ represent the L1 loss and the DINO loss [10], respectively. Following common practices in self-supervised learning [10, 26], sg[$\cdot$] is applied to block the gradients from back-propagating to the observer branch. The weights of the observer branch, including the MLPs, are updated at each training iteration using an EMA of the predictor’s.
Self-supervision for learning powerful representations. When humans learn to interact with the environment, they first develop an understanding of the observed world and objects within, which then aids them in predicting future states of the world and making appropriate responses. Inspired by this process, we introduce an additional self-supervised learning loss on EgoAgent to facilitate learning representative features from egocentric videos from scratch:
@@
Overall objective function. Given an input interleaved token sequence containing $t$ time steps, we have the overall objective function $\mathcal{L}$ defined as:
@@
where $\lambda_{rep}$, $\lambda_{pred}$, and $\lambda_{act}$ are the corresponding loss weights for representing the world, predicting future world states and predicting actions, respectively.